Find the exact value of the volume of a can of tomato paste with a diameter 5 centimeters and height of 8 centimeters. Include correct units with your solution. Use "pi" in place of the pi symbol.
Found 2 solutions by Alan3354, lwsshak3:
Answer by Alan3354(69443)
Answer by lwsshak3(11628)

Find the exact value of the volume of a can of tomato paste with a diameter 5 centimeters and height of 8 centimeters. Include correct units with your solution. Use "pi" in place of the pi symbol.
It might help to visualize volumes as areas moving a certain distace.
In this problem we can calculate the area of the top of the can then move it a distance equal to the height of the can.
area of top of the can = pi*r^2cm^2 =pi*(5/2)^2cm^2=pi(25/4)cm^2
volume = height*area = 8cm*pi(25/4)cm^2=50picm^3
ans:The exact volume of the can is 50picm^3